OpenStudy (anonymous):

MEDAL N FOLLOW 1 QUESTION

OpenStudy (anonymous):

Evan is painting a stool that has a cylindrical seat and four rectangular prism legs. The seat has a diameter of fourteen inches and a height of 4 inches. Each leg is 4 inches by 4 inches by eighteen inches. How much area will Evan have to paint? Use three point one four for pi. Show your work.

OpenStudy (anonymous):

get surface area of cylinder and the legs

OpenStudy (anonymous):

Area of the top of the stool: A = pi*r^2 A = (3.14)*(7)^2 ... The diameter is 14 in, so the radius is 14/2 = 7 in A = 153.86 square inches ----------------------------------- Lateral surface area of the top of the stool A = 2pi*r*h A = 2*(3.14)*7*4 A = 175.84 square inches ------------------------------------- Surface area of one rectangular prism leg (ignoring the top face) SA = 2*(LW+WH+LH) - LW SA = 2*(4*4+4*18+4*18) - 4*4 SA = 2*(4*4+4*18+4*18) - 4*4 SA = 304 There are 4 of these legs. So the total surface area of the legs is 4*304 = 1216 square inches. -------------------------------------- Now add up the area of the top, the sides of the stool, and the legs of the stool to get 153.86 + 175.84 + 304 = 633.7 So Evan will have to paint 633.7 square inches total.

OpenStudy (anonymous):

?/

OpenStudy (anonymous):

its 633.7 Sq in. -_-

OpenStudy (anonymous):

lmao hope i helped

OpenStudy (anonymous):

Marna wants to store one hundred fifty cubic inches of homemade pasta sauce in a food storage container. Should she choose the cylinder with a radius of 3 inches and a height of 5 inches, or the cube with a side length of 6 inches? Explain.

OpenStudy (anonymous):

The formula for the volume of a cylinder is V=πr^2ℓ where r is the radius and l is the side length which measuerments give the larger volume

OpenStudy (anonymous):

pi*r^2*h=pi*9*5=45pi=141.37.. u want 150, so no u won't be able to store

OpenStudy (anonymous):

thats your answer ^

OpenStudy (anonymous):

Celina is buying fabric scraps to make koozies for her glasses. Each glass is cylindrical with a diameter of 4 inches and a height of 8 inches. The koozies will cover only the sides of the glasses. What is the amount of fabric she will use? What does this amount represent? Use three point one four for pi. Show your work.

OpenStudy (anonymous):

make another question and tag me so the mods dont complain plz